A rock is thrown horizontally from the top of a building with a velocity 12.2m/s the rock lands on the ground 103.0m from the edge of the building what is the height of the building? what is the rock's velocity just before hitting the ground?
Answer #1

horizontal velocity of the rock, u = 12.2 m/s

now the distance covered can be found using => R = u*t

we found time taken during the fall is ,,,,, t = R/u = 103 / 12.2 = 8.442 s

height of the buikding can be found using => h = (1/2)*g*t2

we get ...........h = 0.5*9.81*(8.442)^2 = 349.61 m

rock's velocity just before hitting =>

vx = 12.2 m/s

vy = -gt = -9.81 * 8.442 = -82.81 m/s